# Detect, Explain, Escalate: Low-Carbon Dialogue Breakdown Management for LLM-Powered Agents

### v2

@@ V-E Inference and Cost Considerations
For a comprehensive study, we examine a diverse range of LLMs, as listed in Table II. The models are ordered beginning with proprietary models from organizations such as OpenAI (GPT-3.5 and GPT-4[1], Anthropic (Claude-3.5 Haiku and Claude-3.5 Sonnet v2[2]. In contrast, open-source models from Mistral AI (Mixtral 8x7B and Mixtral 8x22B[20], DeepSeek (DeepSeek-R1[21] and Meta (Llama-3.1 8B, Llama-3.3 70B and Llama-3.1 405B[3] offer greater transparency regarding model details, increased flexibility for customization, and are often more cost-effective. For inference, we route queries through the OpenRouter API  [^2] https://openrouter.ai, which provides a unified interface to multiple LLM endpoints. This approach allowed us to evaluate different models under a consistent framework and logging. However, using these hosted models incurs costs per token. In practice, Claude 3.5 Sonnet v2 is the most expensive, roughly on the order of \$5/M input tokens and \$15/M output tokens. We carefully designed prompts to be concise (especially for zero-shot and few-shot setups) to control token usage. Techniques like AR and CL+AR produce longer interactions since the model generates examples as part of its answer, which we note as a trade-off. They may improve reasoning but use more tokens, impacting cost and latency. All models use a temperature of 0. While larger models support longer contexts, a uniform maximum token limit of 2048 was applied across all models to ensure consistent comparison, particularly reflecting the constraints of some smaller models and API limitations for longer AR/CL+AR generations.
@@ VI Results and Discussion
This section presents detailed performance analyses across three datasets: BETOLD (task-oriented dialogues) and DBDC5 (English and Japanese tracks). We evaluate multiple prompting strategies, proprietary and open-source models, and our Dialogue Disruption Monitor. Table II summarizes accuracy, F1 scores for Breakdown (B) and Non-Breakdown (NB) classes, and calibration performance (MSE) where applicable.
TABLE II: Results of proprietary and open-source models on dialogue breakdown datasets. Bold = best, underlined = second‐best.
@@
Accuracy F1(B) F1(NB) Accuracy F1(B) F1(NB) MSE Accuracy F1(B) F1(NB) MSE
Prev. SOTA S2T2 77.9 82.4 76.7 75.4
@@ VI-A Main Findings
State-of-the-Art Results. Our results confirm significant progress beyond earlier benchmarks. On the DBDC5 English track, multiple closed and open-source modern LLMs surpass the prior best (77.9% accuracy from S2T2). For instance, Claude-3.5 Sonnet v2 attains up to 85.5% accuracy (AR prompt), matching Llama-3.3 70B with CL+AR (85.5%). This tie represents the new top performance on DBDC5 English. On the DBDC5 Japanese track, the best outcome of 89.0% accuracy comes from Claude-3.5 Sonnet v2 with CL+AR, slightly above the leading open-source DeepSeek-R1 model at 87.0%. Thus, for Japanese, the Anthropic model retains a slight edge. Nonetheless, the top open-source systems now perform competitively, typically within 1–3 points of the best closed-source counterpart.
Closed-Source Frontier Models. Claude-3.5 (Haiku, Sonnet v2) demonstrate strong classification consistency across both DBDC5 tracks, ranging from 74% to 89% on Japanese and 78% to 85% on English. In particular, Sonnet v2 combined with AR or CL+AR prompts yields top accuracies, for instance, 85.5% on English and 89.0% on Japanese. GPT-4o likewise competes closely, achieving up to 77.7% on BETOLD via 4S prompting and 83.5% on DBDC5 English using a more challenging 2S-Hard strategy. While marginally behind Claude-3.5 Sonnet v2 on Japanese, GPT-4o’s performance remains robust, although it exhibits greater sensitivity to variations in prompt style.
In contrast, GPT-3.5 Turbo underperforms substantially on BETOLD, with accuracy ranging from 41% to 64.5%. It tends to misclassify borderline “near-breakdown” utterances or produce imbalanced predictions. A plausible explanation is that GPT-3.5 Turbo is more sensitive to how examples are presented; certain prompt structures lead to skewed confidence or confusion in distinguishing near-breakdown from non-breakdown scenarios.
@@
Dialogue Disruption Monitor. Despite having fewer parameters, our fine-tuned Llama-3.1 8B model achieves competitive performance. On the DBDC5 English dataset, it attains an accuracy of 81.5%, surpassing several larger models. For the DBDC5 Japanese dataset, the model achieves a moderate yet stable accuracy of 67.9%, with balanced F1 scores across classes. Our DBDC5-fine-tuned monitor, when evaluated on the BETOLD dataset (on which it was not trained), achieved 67.2% accuracy. This represents a 7% absolute improvement over the base Llama-3.1 8B model’s zero-shot performance (60.2%) on BETOLD, indicating effective transfer of breakdown detection capabilities to a new dataset and format. On DBDC5 English, the model’s performance is comparable to that of significantly larger Llama variants while demonstrating better calibration scores. However, performance on the Japanese dataset remains limited due to insufficient Japanese training data. Future research could benefit from utilizing models such as Llama-3.1 Swallow 8B, which has undergone additional training on 200 billion tokens derived from the extensive Japanese web corpus (Swallow Corpus Version 2), as well as Japanese and English Wikipedia [44].
@@ VI-B Impact of Prompting Strategies
Few-Shot Prompting. Few-shot prompting consistently outperforms ZS and CoT approaches across datasets. For instance, DeepSeek-R1 rises from 81.1% (ZS) to 83.0% (4S). The “Hard” exemplars comprised of borderline dialogues also yield stronger calibration (lower MSE). 2S-Hard achieves substantial calibration improvements, with Llama-3.3 70B reaching an MSE of 4.0 on DBDC5 English, the lowest among all models. Similarly, GPT-4o’s accuracy improves from 81.4% (ZS) to 83.5% (2S-Hard), affirming the effectiveness of providing challenging examples to refine model uncertainty estimations more effectively than simpler or even more examples.
@@
Challenges with AR and CL+AR on BETOLD. While AR and CL+AR techniques generally improved performance on natural dialogue datasets (such as the DBDC5 tracks), their effectiveness significantly declined when applied to BETOLD. Results for BETOLD (highlighted in grey in Table II) consistently showed lower performance. We hypothesize this reduction is primarily due to the nature of BETOLD itself, which represents dialogues through structured intents and entities rather than complete natural language utterances. This mismatch likely hinders the models’ generalization capabilities since they were predominantly trained on conventional dialogue. Through comprehensive error analysis, we observed that models frequently did not fully adhere to the provided instructions. Common issues included either failing to generate any analogous examples, opting instead to solve the original dialogue directly, or producing overly brief analogous summaries that poorly aligned with the dialogue under evaluation. As demonstrated in Figure 4, even frontier models like GPT-4o exhibited challenges in accurately following instructions. This problem is exacerbated by the relatively longer dialogues in BETOLD (averaging 20-30 turns) compared to the shorter dialogues in DBDC5 English and Japanese tracks (typically 10-20 turns). Providing three analogous examples plus restating the original conversation often exhausted the available token limit, as depicted in Figure 3. Moreover, smaller model variants such as Llama-3.1 8B, Mixtral 8x7B, and Claude-3.5 Haiku struggled to generate relevant analogies compared to their larger counterparts.
@@ VI-C Calibration and Confidence Analysis
Superior calibration (low MSE) indicates alignment between model confidence and human correctness. Llama-3.3 70B reaches the best DBDC5 English calibration (4.0, 2S-Hard), while Llama-3.1 405B achieves 4.6 on Japanese (2S-Hard). Among closed-source, Claude-3.5 Haiku can hit 4.1 MSE on DBDC5 English (2S-Hard). Meanwhile, our fine-tuned 8B Dialogue Disruption Monitor yields an MSE of 4.9. While slightly higher, this result still demonstrates that moderate‑scale, task‑specific training can produce stable confidence alignment.
VI-D Practical Costs and Sustainability
@@
Resource efficiency is a crucial factor influencing the practical deployment of language models. Advanced models such as Claude-3.5 Sonnet v2 (400B parameters), DeepSeek-R1 (671B parameters), and GPT-4 (estimated 1.7 trillion parameters[45]) demand substantial monetary costs, experience network-induced latency, and have a higher carbon footprint per query due to their considerable size and reliance on data-center computation. For example, based on cloud-provider pricing from Amazon Web Services (AWS), querying Llama-3.1 70B is approximately 3.3 times more expensive than its 8B-parameter counterpart, while Llama-3.1 405B incurs a cost approximately 10.3 times higher [46]. Additionally, using longer prompts (such as AR or CL+AR) significantly escalates costs due to the increased token usage per query.
In contrast, our fine-tuned Llama-3.1 8B model operates efficiently on a single A100 GPU, resulting in minimal incremental cost per query and low inference latency (each dialogue inference completes in under half a second). This efficiency is advantageous for deployments requiring real-time detection of dialogue breakdowns for every utterance. Assuming only 10% of utterances lead to dialogue breakdown, our practical deployment architecture is estimated to reduce costs by approximately 54% compared to employing Llama-3.1 405B for each turn (see Appendix A for detailed cost calculations). Such selective invocation lowers operational costs and aligns with sustainability objectives. Enterprise-scale LLM deployments like ChatGPT consume around 1,058.5 GWh annually, equivalent to the yearly electricity consumption of approximately 100,810 U.S. households [47]. By limiting resource-intensive model invocations to instances of suspected critical dialogue breakdowns or high uncertainty, our approach effectively balances performance, cost-efficiency, and sustainability.
@@ VII Conclusion
Research and development in conversational AI robustness is rapidly advancing on multiple fronts. Larger and better-trained LLMs provide a stronger base, specialized benchmarks like DBDC5 and BETOLD drive progress in error detection, and techniques such as analogical reasoning, curriculum learning, and chain-of-thought help models use their internal knowledge more effectively. By combining high-performing models with these strategies, the latest conversational agents are increasingly adept at sustaining error‑free dialogues; however, evaluations show that a gap remains between current systems and human‑level communication when handling the full complexity of conversation. While advanced prompting (CoT, AR) can improve performance for some high-parameter models, these gains are inconsistent in specialized domains. Short, curated exemplars typically offer a good trade-off between performance and token usage. The cost and energy savings are considerable, especially in frequent-turn applications. Eliciting numeric confidence and justifications can mitigate overconfidence and improve trustworthiness. Our experiments reveal that carefully structured prompts, where the model justifies and then decides, yield lower calibration error. For high-stakes, high-volume systems, coupling a fast, fine-tuned breakdown detector with on-demand escalation to a large frontier LLM can ensure both reliability and sustainability. While closed-source models still lead in absolute accuracy, the performance gap is closing, especially once small open-source models are carefully tuned. Future work should focus on systematically combining these approaches to achieve robust, interpretable breakdown detection at scale.
@@ VIII Limitations
Although our approach performs competitively on both English and Japanese DBDC5 benchmarks, the smaller Llama-3.1 8B model shows limited Japanese coverage, suggesting stronger multilingual pretraining is needed. Moreover, real-world dialogues (involving code-switching, adversarial inputs, or varied domains) may demand specialized adaptations beyond our well-defined benchmarks, while prompt engineering remains model- and dataset-specific, increasing resource demands. The opaque nature of LLMs, even under chain-of-thought or analogical prompting, leaves internal reasoning partially hidden; consequently, escalating errors to more costly, higher-capacity models adds latency, energy use, and expense. Finally, simpler binary breakdown labels may miss borderline cases, pointing to a need for finer-grained schemes and more resource-aware architectures for robust deployment in diverse, high-stakes environments. Nonetheless, these constraints do not undermine the contributions of this work; we leave their resolution to future research.

### v3

@@ V-E Inference and Cost Considerations
We evaluate a broad set of LLMs, summarized in Table II. Proprietary models include OpenAI’s GPT-3.5 Turbo and GPT-4o [openai_gpt-4_2023], and Anthropic’s Claude-3.5 Haiku and Claude-3.5 Sonnet [ClaudeModelCard]. Open-source models include Mistral’s Mixtral 8x7B and Mixtral 8x22B [jiang_mixtral_2024], DeepSeek’s DeepSeek-R1 [deepseek], and Meta’s Llama-3.1 8B, Llama-3.3 70B, and Llama-3.1 405B [llama]. Inference is run via the OpenRouter API [^2] https://openrouter.ai with a temperature of 0. To ensure comparability across models, we cap total tokens at 2,048 per query, which also reflects practical constraints for longer AR/CL+AR prompts.
Advanced prompting strategies such as AR and CL+AR require additional tokens, since models must generate analogous examples in context. We note this increased token usage as a trade-off: these strategies may improve reasoning at higher cost and latency.
@@ VI Results and Discussion
We now analyze performance across BETOLD and the English and Japanese tracks of DBDC5. We compare prompting strategies, model families, and our fine-tuned Dialogue Disruption Monitor. Table II reports accuracy, class-specific F1 scores, and calibration (MSE) where applicable.
TABLE II: Results of proprietary and open-source models. Note: MSE values are scaled by 100 ($10^{-2}$). Bold is best, underlined is second-best. Rows highlighted in grey indicate the BETOLD subset on specific prompting techniques.
@@
Acc. F1B F1NB Acc. F1B F1NB MSE Acc. F1B F1NB MSE
Previous SOTA BERT+SSMBA 73.9 78.2
S2T2 77.9 82.4 76.7 75.4
@@ VI-A Main Findings
State-of-the-Art Results. On DBDC5 English, multiple modern LLMs surpass the prior best accuracy of 77.9% from S2T2. Claude-3.5 Sonnet and Llama-3.3 70B both reach 85.5% accuracy under AR and CL+AR prompting, respectively, achieving state-of-the-art performance. On DBDC5 Japanese, Claude-3.5 Sonnet with CL+AR attains 89.0% accuracy, slightly ahead of the strongest open-source model, DeepSeek-R1, at 87.0%. Overall, open-source models are now within 1–3 points of leading closed-source systems on these benchmarks.
Closed-Source Frontier Models. Claude-3.5 (Haiku, Sonnet) demonstrate strong classification consistency across both DBDC5 tracks, ranging from 74% to 89% on Japanese and 78% to 85% on English. In particular, Sonnet combined with AR or CL+AR prompts yields top accuracies, for instance, 85.5% on English and 89.0% on Japanese. GPT-4o likewise competes closely, achieving up to 77.7% on BETOLD via 4S prompting and 83.5% on DBDC5 English using a more challenging 2S-Hard strategy. While marginally behind Claude-3.5 Sonnet on Japanese, GPT-4o’s performance remains robust, although it exhibits greater sensitivity to variations in prompt style.
In contrast, GPT-3.5 Turbo underperforms substantially on BETOLD, with accuracy ranging from 41% to 64.5%. It tends to misclassify borderline “near-breakdown” utterances or produce imbalanced predictions. Error analysis suggests that GPT-3.5 Turbo is more sensitive to how examples are presented; certain prompt structures lead to skewed confidence or confusion in distinguishing near-breakdown from non-breakdown scenarios.
@@
Dialogue Disruption Monitor. Despite its size, our fine-tuned Llama-3.1 8B monitor achieves 81.5% accuracy on DBDC5 English, exceeding several larger models and approaching Llama-3.3 70B. On DBDC5 Japanese, it reaches 67.9% accuracy with balanced F1 scores, reflecting limited Japanese coverage in pretraining. On BETOLD, the monitor achieves 67.2% accuracy, a 7-point absolute improvement over the zero-shot Llama-3.1 8B baseline (60.2%). This demonstrates cross-dataset transfer of breakdown detection behaviour. On DBDC5 English, the monitor’s calibration (MSE = 4.9) is competitive with much larger models. To improve Japanese performance further, future work could explore models such as Llama-3.1 Swallow 8B, which are pre-trained on large-scale Japanese corpora [swallow:COLM2024].
@@ VI-B Impact of Prompting Strategies
Few-Shot Prompting. Few-shot prompting consistently outperforms ZS and CoT approaches across datasets. For instance, DeepSeek-R1 rises from 81.1% (ZS) to 83.0% (4S). The “Hard” exemplars comprised of borderline dialogues, also yield stronger calibration (lower MSE). 2S-Hard achieves substantial calibration improvements, with Llama-3.3 70B reaching an MSE of 4.0 on DBDC5 English, the lowest among all models. Similarly, GPT-4o’s accuracy improves from 81.4% (ZS) to 83.5% (2S-Hard), affirming the effectiveness of providing challenging examples to refine model uncertainty estimations more effectively than simpler or even more examples.
@@
AR and CL+AR on BETOLD. While AR and CL+AR techniques improved performance on natural dialogue datasets (DBDC5 tracks), their effectiveness declined significantly on BETOLD. Results (highlighted in grey in Table II) consistently showed lower performance. This degradation stems from BETOLD’s representation of dialogues through structured intents and entities rather than natural language utterances. This mismatch undermines generalization, as models were predominantly trained on conventional dialogue data. Error analysis revealed frequent instruction-following failures. Models either generated no analogous examples (instead directly solving the original dialogue) or produced overly brief analogies with weak alignment to the target dialogue. Even GPT-4o struggled with reliable compliance (Figure 5). BETOLD’s longer dialogues (20–30 turns vs. 10–20 turns in DBDC5) exacerbate these issues. Three analogous examples plus the original conversation frequently exhausted the token budget (Figure 4). Smaller variants (Llama-3.1 8B, Mixtral 8x7B, Claude-3.5 Haiku) showed poor analogy generation, likely due to their reduced capacity and limited training.
@@ VI-C Calibration and Confidence Analysis
Low MSE indicates good alignment between model confidence and human agreement. On DBDC5 English, Llama-3.3 70B with 2S-Hard achieves the best calibration (MSE = 4.0), while Llama-3.1 405B with 2S-Hard obtains the best calibration on Japanese (MSE = 4.6). Among proprietary models, Claude-3.5 Haiku reaches MSE = 4.1 on English under 2S-Hard. Our 8B monitor’s MSE of 4.9 demonstrates that moderate-scale, task-specific fine-tuning can yield calibration quality close to much larger models.
VI-D Sensitivity Analysis
We analyze how the escalation threshold $T$ trades off safety and cost. For each dataset, we vary $T$ (minimum confidence required to accept a Non-Breakdown decision without escalation) and report safety (recall for the Breakdown class) and cost (escalation rate, i.e., the fraction of turns escalated).
Decisive Regime ($T\leq 0.5$). Across DBDC5 English, DBDC5 Japanese, and BETOLD, the curves are largely flat for $T\in[0,0.5]$, indicating that the monitor is confident when predicting Non-Breakdown. In this regime the system behaves like a near-binary classifier with strong safety at low manual tuning effort. For example, on DBDC5 English, we obtain $\approx$88% recall while escalating only $\approx$70% of turns, about a 30% reduction in compute relative to full escalation.
Calibration Regime ($T>0.5$). Beyond $T=0.5$, the monitor begins escalating low-confidence non-breakdown predictions, yielding coupled increases in safety and cost. This reflects an effective uncertainty safety net: raising $T$ recovers borderline breakdowns that would otherwise be missed. On the out-of-domain BETOLD set, increasing $T$ from 0.5 to 0.8 boosts safety from roughly 70% to $90\%$, at the expense of higher escalation.
Operational Guidance. The architecture exposes a simple knob for budget–safety trade-offs. A default of $T=0.5$ yields substantial efficiency gains for standard deployments. For high-stakes domains where missed breakdowns are unacceptable, $T$ can be increased to push breakdown recall closer to 100% while still avoiding full reliance on the largest model.
(a) DBDC5 English: High baseline safety with $\sim$30% fewer escalations.
(b) DBDC5 Japanese: Similar profile, indicating multilingual stability.
(c) BETOLD: Lower due to domain shift; tuning $T$ pushes safety >>90%.
Figure 3: Sensitivity to the escalation threshold. As $T$ increases, safety (breakdown recall) and cost (escalation rate) trace a Pareto-like frontier. The flat segment for $T\leq 0.5$ reflects decisive Non-Breakdown predictions; the rising segment for $T>0.5$ shows that computation can be traded for additional safety.
VI-E End-to-End Repair Capabilities
To address the need for closed-loop remediation, we evaluated the effectiveness of the ”Escalate” module in repairing detected breakdowns. We sampled 100 correctly detected breakdowns from the DBDC5 English test set. We passed the dialogue history and the monitor’s generated justification ($\hat{j}_{i}$) to the Superior Model (Claude-3.5 Sonnet) with instructions to repair the response.
Automated evaluation using an LLM-as-a-judge (configured to assess response validity) confirmed that the pipeline successfully resolved the breakdown in 97% of cases [Gu2024A]. Qualitative analysis reveals that the superior model effectively utilizes the monitor’s diagnostic context to generate coherent replacements. For example, when the assistant hallucinated a contradiction (”I do not play video games. I like to play video games”), the repair module used the monitor’s flag to generate a consistent, context-aware correction (”I do enjoy playing video games in my free time…”). This confirms that the pipeline effectively converts detection signals into actionable repairs.
VI-F Resource Efficiency and Sustainability
@@
Resource efficiency is a crucial factor influencing the practical deployment of language models. Advanced models such as Claude-3.5 Sonnet (400B parameters), DeepSeek-R1 (671B parameters), and GPT-4 (estimated 1.7 trillion parameters[epochFrontierLanguage]) demand substantial monetary costs, experience network-induced latency, and have a higher carbon footprint per query due to their considerable size and reliance on data-centre computation. For example, based on cloud-provider pricing from Amazon Web Services (AWS), querying Llama-3.1 70B is approximately 3.3 times more expensive than its 8B-parameter counterpart, while Llama-3.1 405B incurs a cost approximately 10.3 times higher [amazonbedrockpricing]. Note, this also does not consider the latency, which in practice also increases substantially with model size, network routing complexity, and prompt length. Additionally, using longer prompts (such as AR or CL+AR) significantly escalates costs due to the increased token usage per query.
In contrast, our fine-tuned Llama-3.1 8B monitor runs efficiently on a single A100 GPU, with latency under half a second per dialogue turn, making it suitable for per-turn monitoring. In our cost analysis (Appendix A), we estimate that a hierarchical deployment combining a 70B assistant, 405B escalations on 10% of turns, and an 8B monitor achieves a 54% cost reduction relative to running the entire dialogue on Llama-3.1 405B. Given that large deployments such as ChatGPT may consume on the order of 1,058.5 GWh annually [bestbrokersAIsPower], selectively invoking large models only when necessary can meaningfully reduce cost and environmental impact.
VI-G Relevance of Breakdown Datasets to Modern LLMs
Although DBDC5 was developed for chat-oriented dialogue systems rather than modern LLM deployments, its breakdown labels and error categories continue to capture core failure modes that remain prevalent in current deployments [Higashinaka2020DBDC5]. DBDC5 explicitly models phenomena such as contradiction, wrong or missing information, ignored questions, and topic-transition errors, which map closely onto “intrinsic hallucinations” (conflicts with source context) and broader faithfulness errors in recent hallucination taxonomies [Huang2025Taxonomy, Higashinaka2020DBDC5]. Consequently, detecting these breakdown types remains a useful proxy for assessing LLM reliability on fundamental conversational errors, even as underlying architectures and training regimes evolve.
@@ VII Conclusion
Robustness in conversational AI is advancing along several dimensions: larger and better-trained LLMs, more targeted benchmarks such as DBDC5 and BETOLD, and reasoning techniques such as chain-of-thought, analogical prompting, and curriculum learning. Our work combines these elements into a practical framework for dialogue breakdown management. Empirically, we show that modern LLMs, both proprietary and open-source, achieve state-of-the-art performance on dialogue breakdown detection benchmarks. However, a gap remains to human-level performance, especially in handling complex and ambiguous conversations. Advanced prompting (e.g., CoT, AR) can yield further gains for some high-capacity models, but improvements are not uniform across datasets. Short, well-chosen exemplars often strike a better balance between performance and token usage. We find that eliciting both justifications and numeric confidence can reduce overconfidence and improve calibration, particularly when the model is prompted to justify before deciding. For high-volume, high-stakes applications, coupling a fast, fine-tuned monitor with on-demand escalation to a frontier LLM provides a practical path to reliability and sustainability. While closed-source models currently lead in absolute accuracy, careful tuning of small open-source models narrows the gap. Future work should explore more fine-grained breakdown taxonomies, richer multimodal settings, and tighter integration of monitoring signals into policy-level decisions, with the goal of building robust, interpretable, and resource-aware conversational systems at scale.
@@ VIII Limitations
Although our approach performs competitively on both English and Japanese DBDC5 benchmarks, the smaller Llama-3.1 8B model shows limited Japanese coverage, suggesting that stronger multilingual pretraining or dedicated Japanese adaptation (e.g., using Swallow [swallow:COLM2024]) is needed. Moreover, real-world dialogues involving code-switching, adversarial inputs, or highly specialized domains may demand additional adaptations beyond the benchmarks we evaluate. Second, prompt engineering remains model- and dataset-specific, and methods such as AR and CL+AR can be brittle and expensive. Moreover, LLMs remain partially opaque: even with chain-of-thought or analogical prompting, internal reasoning processes are not fully observable. Escalating to larger models introduces latency and energy overhead, which may be problematic in ultra-low-latency or resource-constrained settings. Finally, our binary breakdown labels elide nuanced distinctions between mild and severe breakdowns and do not explicitly capture repair quality or user satisfaction. More expressive annotation schemes and joint modelling of detection and repair could enable finer-grained control in deployment. We leave these directions to future work.